OpenStudy (anonymous):

what is 44 conected with NaOH?

OpenStudy (mos1635):

44?...are you sure??

OpenStudy (anonymous):

yes, i found it in equals \[ppm CO2 with NaOH= \frac{ 1000 }{ ml sample } \times ml titration \times normalitate NaOH \times 44 what is 44??\]

OpenStudy (aaronq):

lol what?

OpenStudy (mos1635):

looks like Molecular weith of CO2